OpenStudy (anonymous):

Medal To discourage guessing on a multiple-choice exam, an instructor graded the test by giving 3 points for a correct answer, 0 points for an answer left blank, and -1 points for an incorrect answer. How many points did a student score who answered 36 questions correctly, answered 8 questions incorrectly, and left 6 questions blank? ? points

OpenStudy (anonymous):

Well, the total number of points would be the correct answers *3 plus incorrect answers *-1. So it would be P=3*36-1*8=108-8=100 points.

OpenStudy (anonymous):

The maximum number of points would be total number of questions *3. So Max=3*(36+8+6)=3*50=150 points possible.

OpenStudy (anonymous):

Therefore, this student got 100 out of 150 points, or a 66%.

OpenStudy (anonymous):

so the answer is 150 points?

OpenStudy (anonymous):

No, the answer is 100 points out of a possible 150 points,

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

is not the answer 96

OpenStudy (anonymous):

Then you must have stated the question wrong, because 36*3=108 and the student got -8 from incorrect answers, therefore they got 100 points.

OpenStudy (anonymous):

what about the questions they left blank

OpenStudy (welshfella):

they get 0 points for a question left blank

OpenStudy (anonymous):

so the answer is 100 points?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

thank you